A stellar object is emitting radiation at 1620 nm. A detector is capturing 9 107 photons per second at this wavelength. h = 6.63 x 10-34 J-s. c = 2.998 x 108 m/s.

(a) What is the total energy of the photons detected in one second?
J/s

(b) What is the total energy of the photons detected in one hour?
J/hr
Answer #1

(a): energy of one photon = h* c / λ

In one second 9* 107 photons are emited.

So the energy on 9107 photons will be = 9* 107* h* c / λ

= 9* 107* 6.63 x 10-34* 2.998 x 108 / (1620* 10-9)

= 1.1036* 10-11 J/Sec

(b): in 1 hour the total energy emitted = 3600* energy emitted in one second

= 3600* 1.1036* 10-11

= 3.9729* 10-8 J